3 equal coils of inductive reactance 30 ohms and resistance 40 ohms are connected in Star to a 3 phase supply of line voltage 400 volts. CALCULATE the line Current and the Total Power


1
Expert's answer
2021-03-04T06:03:12-0500

"Z_p=\sqrt{30^2+40^2}=50\ (\Omega)"


"V_p=400/\sqrt{3}=231\ (V)"


"I_L=I_p=V_p/Z_p=231/50=4.62\ (A)" . Answer


"S=\sqrt3V_LI_L=\sqrt3\cdot400\cdot4.62=3201\ (W)" . Answer